RATE BACK ASAP!!!

pls help me answer BOTH parts of this question. Multiple choice answers. I would really appreciate it, as I need help on BOTH parts. Thank you so much in advance.

Find VO as a function of the independent sources. In essence, (Zal,Va)-ala + ßVa find α and 3. 552 352 What are α and β α-1.6.9-0.2 ο α-2.9-0.2 If Vout16V when all the sources are on, what would it equal if the 16V source were turned off e. shorted)? Hint: Do you need to consider the ontribution of all the sources? 24V Vout 42 45 2 2t? 41 ? 4A 2A 16V 2Ω What is Vo when the 16V souroe is turned off? 20V 12V ●AV -4V 16.25V 15.75V
Answer #1

и 5 - 3 ~ фото то Kut for loop @ - - Vi2 + 3), +si, +20 .-) = 0 oi, - 2 = Ч, 2 - 2 + 2) Э 10 .: М. 2 x y -) - 2 х (- 42-211]Tuntem M Live Kel at - U + 2 V2 – V; =-16-0 super node Equation 20,- 2.0+ 30, -30,+30373 02 + 603 = 24 → 541-602 + 9 U₂ = 64